### v1

# Overcoming Knowledge Barriers: Online Imitation Learning from Observation with Pretrained World Models

Abstract
Incorporating the successful paradigm of pretraining and finetuning from Computer Vision and Natural Language Processing into decision-making has become increasingly popular in recent years. In this paper, we study Imitation Learning from Observation with pretrained models and find existing approaches such as BCO and AIME face knowledge barriers, specifically the Embodiment Knowledge Barrier (EKB) and the Demonstration Knowledge Barrier (DKB), greatly limiting their performance. The EKB arises when pretrained models lack knowledge about unseen observations, leading to errors in action inference. The DKB results from policies trained on limited demonstrations, hindering adaptability to diverse scenarios. We thoroughly analyse the underlying mechanism of these barriers and propose AIME-v2 upon AIME as a solution. AIME-v2 uses online interactions with data-driven regulariser to alleviate the EKB and mitigates the DKB by introducing a surrogate reward function to enhance policy training. Experimental results on tasks from the DeepMind Control Suite and Meta-World benchmarks demonstrate the effectiveness of these modifications in improving both sample-efficiency and converged performance. The study contributes valuable insights into resolving knowledge barriers for enhanced decision-making in pretraining-based approaches. Code will be available at https://github.com/argmax-ai/aime-v2.
1 Introduction
In recent years, we are going through a paradigm shift from learning from scratch to pretraining and finetuning in Computer Vision (CV) [25, 47, 26] and Natural Language Processing (NLP) [12, 46, 43, 59, 60] fields due to the increasing availability of foundation models [6] and ever growing datasets. However, how to adapt this new paradigm into decision making is still unclear, i.e. what type of models we need to pretrain and how these models can be adapted to solve downstream tasks. Recent works [65, 11, 52, 48, 23] showcase that one can pretrain a shared latent space world model, enabling successful and efficient transfer to new tasks with either reinforcement learning [52, 48, 23] or Imitation Learning from Observation (ILfO) [65, 11]. Arguably, ILfO [57, 58, 3, 65, 11, 37], especially from videos [3, 65, 37, 11], is a more promising approach in this new paradigm since it does not require a handcrafted reward function which is hard to define for many real-world tasks.
However for ILfO with pretrained models, we identify two barriers, which limit the effectiveness of employing pretrained models, the Embodiment Knowledge Barrier (EKB) and the Demonstration Knowledge Barrier (DKB). The EKB describes the shortcomings of a pretrained model when confronted with novel observations and actions beyond its training experience. The DKB describes the known problem of generalisation from a limited number of expert demonstrations in imitation learning [27]. We find current approaches like BCO [57] and AIME [65] typically suffer from these two knowledge barriers. First, these algorithms depend on the pretrained model to successfully infer missing actions from observation sequences. Thus, when the model has not seen a specific observation before, it may not contain sufficient knowledge about the embodiment to infer the correct action. Second, policy optimisation is only guided by a limited demonstrations, which may not provide sufficient coverage of the space and lead to a brittle policy working well in some scenarios but not others.
Figure 1: Main idea of this paper. On the left, we extend the plots from Figure 4 in Zhang et al. [65] by adding the MBBC variant and draw the BCO(0) and AIME separately. The purple region between the oracle version and the expert is the DKB while the orange region between the algorithm and its oracle version represents the EKB. On the right, we present the solutions proposed in this paper to solve the two barriers. The blue parts represent the original version of the algorithms that suffer from the knowledge barriers. Orange parts demonstrate the solution for EKB, where the agent is allowed to interact with the environment and use $D_{\mathrm{online}}$ together with $D_{\mathrm{body}}$ to update the world model. Purple parts show the solution for DKB, where a VIPER model is pretrained and used to label the online dataset and then used as a RL signal for policy learning.
To better showcase the two barriers, in Figure 1 we extended Figure 4 from Zhang et al. [65] by adding a model-based BC (MBBC) variant where the policy is learned by running behaviour cloning on the learned latent space of the world model. BCO(0) [57] and AIME both pretrain a model from a large embodiment dataset and use that to infer the actions for the observation-only demonstrations. The oracle versions remove the need to infer the missing actions, thus removing the EKB. As we can see from the figure, the two algorithms are always upper-bounded by the corresponding oracle version, and the difference between them represents the EKB. On the other hand, even if given the true actions of the expert, imitation performance may still be impacted by a limited number of demonstrations providing insufficient coverage of the state space. Thus, the difference between the oracle version and the expert performance represents the DKB.
In this paper, we study how to resolve these barriers to improve the performance of ILfO approaches, in particular of AIME proposed by Zhang et al. [65]. For the EKB, we extend the setting from offline to online by allowing the agent to further interact with the environment to gather more data to train the world model. While for the DKB, we introduce a surrogate reward function to allow the policy to essentially train on more data. We demonstrate that the proposed modifications significantly improve the performance on $9$ tasks in DeepMind Control Suite (DMC) [61] and $6$ tasks in Meta-World [64].
2 Preliminary
We mostly follow the problem setup as described in Zhang et al. [65]. We consider a POMDP problem defined by the tuple $\{S,A,T,R,O,\Omega\}$, where $S$ is the state space, $A$ is the action space, $T:S\times A\rightarrow S$ is the dynamic function, $R:S\rightarrow\mathbb{R}$ is the reward function, $O$ is the observation space, which is image in this paper, and $\Omega:S\rightarrow O$ is the emission function. The goal is to find a policy $\pi:S\rightarrow A$ which maximises the accumulated reward, i.e. $\sum_{t}r_{t}$.
We presume the existence of three datasets of the same embodiment available to our agent. The embodiment dataset $D_{\mathrm{body}}$ contains trajectories $\{o_{0},a_{0},o_{1},a_{1}\dots\}$ that represent past experiences of interacting with the environment. This dataset provides information about the embodiment for the algorithm to learn a world model. In addition, we also allow the agent to interact with the environment to collect new data in a replay buffer $D_{\mathrm{online}}$. Note that, although the simulator will give us the reward information, the agent is not allowed to use them, and we only use the reward for evaluation purpose. The demonstration dataset $D_{\mathrm{demo}}$ contains a few expert trajectories $\{o_{0},o_{1},o_{2}\dots\}$ of the embodiment solving a certain task defined by $R_{\mathrm{demo}}$. The crucial difference between this dataset and the other two datasets is that the actions are not provided anymore since they are not observable from a third-person perspective. The goal of our agent is to learn a policy $\pi$ from $D_{\mathrm{demo}}$ which can solve the task defined by $R_{\mathrm{demo}}$ as well as the expert who generated $D_{\mathrm{demo}}$.
2.1 World Models
A World Model [18] is a generative model which models a probability distribution over sequences of observations, i.e. $p(o_{1:T})$. The model can be either unconditioned or conditioned on other factors, such as previous observations or actions. When the actions taken are known, they can be considered as the condition, i.e.  $p(o_{1:T}|a_{0:T-1})$, and the model is called embodied [65].
In this paper, we consider variational latent world models where the observation is governed by a Markovian hidden state. In the literature, this type of model is also referred to as a State-Space Model (SSM) [30, 20, 19, 5, 33]. Such a variational latent world model involves four components, namely
encoder $z_{t}=f_{\phi}(o_{t}),$ posterior $s_{t}\sim q_{\phi}(s_{t}|s_{t-1},a_{t-1},z_{t}),$
decoder $o_{t}\sim p_{\theta}(o_{t}|s_{t}),$ prior $s_{t}\sim p_{\theta}(s_{t}|s_{t-1},a_{t-1}).$
$f_{\phi}(o_{t})$ is the encoder to extract the features from the observation; $q_{\phi}(s_{t}|s_{t-1},a_{t-1},z_{t})$ and $p_{\theta}(s_{t}|s_{t-1},a_{t-1})$ are the posterior and the prior of the latent state variable; while $p_{\theta}(o_{t}|s_{t})$ is the decoder that decodes the observation distribution from the state. $\phi$ and $\theta$ represent the parameters of the inference model and the generative model respectively.
Typically, the model is trained by maximising the Evidence Lower Bound (ELBO) which is a lower bound of the log-likelihood, or evidence, of the observation sequence, i.e. $\log p_{\theta}(o_{1:T}|a_{0:T-1})$. Given a sequence of observations, actions, and states, the objective function can be computed as
ELBO $=\sum_{t=1}^{T}J_{t}^{\mathrm{rec}}-J_{t}^{\mathrm{KL}}=\sum_{t=1}^{T}\log p_{\theta}(o_{t}|s_{t})-D_{\text{KL}}[q_{\phi}||p_{\theta}].$ (1)
The objective function is composed of two terms: the first term $J^{\mathrm{rec}}$ is the likelihood of the observation under the inferred state, which is usually called the reconstruction loss; while the second term $J^{\mathrm{KL}}$ is the KL divergence between the posterior and the prior distributions of the latent state. To compute the objective function, we use the re-parameterisation trick [32, 50] to autoregressively sample the inferred states from the observation and action sequence.
In summary, a world model is trained by solving the optimisation problem as
$\phi^{*},\theta^{*}=\mathop{\arg\!\max}_{\phi,\theta}\mathbb{E}_{\{o,a\}\sim D_{\mathrm{body}},s\sim q_{\phi}}[\text{ELBO}].$ (2)
2.2 AIME
AIME is a recently proposed algorithm that uses a pretrained world model to solve ILfO in an offline setting. Specifically, it uses the pretrained world model as an implicit inference model by solving for the best action sequence that makes the demonstration most likely under the trained world model. The imitation can be done jointly with the action inference using amortised inference and the re-parameterisation trick by solving the following optimisation problem
$\psi^{*}=\mathop{\arg\!\max}_{\psi}\mathbb{E}_{o\sim D_{\mathrm{demo}},s\sim q_{\phi^{*},\theta^{*}},a\sim\pi_{\psi}}[\text{ELBO}],$ (3)
where $\psi$ is the parameter for policy $\pi_{\psi}(a_{t}|s_{t})$. The resulting objective is very similar to Equation 2, with a subtle difference of the sampling path. That is in the new objective, only the observations are sampled from the dataset and both states and actions are sampled iteratively from the learned model and the policy, respectively.
3 Methodology
In this section we will make thorough analysis separately on the EKB and DKB especially upon the AIME algorithm as the case study. Based on the analysis we introduce the solution for each knowledge barriers and combine them as AIME-v2. We depict the general framework of the solutions on Figure 1 and show the pseudocode of AIME-v2 in Algorithm 1 in Appendix A.
3.1 Resolving the EKB
The most natural way to solve the EKB is to allow the agent to further interact with the environment. By allowing the agent to interact with the environment, the new experiences can minimise the error in the pretrained model in proximity of the policy $\pi_{\psi}$ and gain more embodiment knowledge relevant for the task at hand. Torabi et al. [57] proposed a modified version of BCO(0) called BCO($\alpha$) which introduced such an interaction phase. However, from their and our empirical results, it did not resolve the EKB since there remains a gap between BCO($\alpha$) and the BC oracle when the environment is complex. In fact, as we will show in the following, the basic idea of adding online interactions is not straightforward to successfully implement in practice.
As shown in recent works in Offline RL, continuing training an actor-critic from the offline phase in the online phase requires certain measures to combat the shift of objective [34, 4, 42]. A similar story also applies when extending AIME from purely offline to online. The most dominant problem we found is overfitting to the newly collected dataset.
As the training progresses iteratively between data collection, model training and policy training, in the early phase of training, there are only a few new trajectories available for training the model. Because the world model is highly expressive, it may overly favor similar trajectories in the new data, leading to a high ELBO. Normally, this may not be a big problem since, eventually, more and more data will be collected to combat this overfitting. But since AIME also depends on the ELBO to train the policy, it quickly causes the policy training to diverge.
In order to address the overfitting issue, we need a regulariser for model learning. Instead of designing ad-hoc methods to regularise the model in the parameter space, we adopt a data-driven approach. From model’s perspective, the overfitting is caused by a sudden shift of the training data from a large and diverse pretraining dataset to a small and narrow replay buffer. So one way to make the shift not as sudden is to append the pretraining dataset to the replay buffer, so that the distribution of the training data will change smoothly. However, this causes data efficiency problems since the replay buffer is relatively small compared to the pretraining dataset so that uniformly sampling from them together limits utilisation rate of the new data. Instead, we consider sampling separately from both datasets. Concretely, we modify the objective in Equation 2 to
$\phi^{*},\theta^{*}=\mathop{\arg\!\max}_{\phi,\theta}\>$ $\alpha\mathbb{E}_{\{o,a\}\sim D_{\mathrm{body}},s\sim q_{\phi}}[\text{ELBO}]+(1-\alpha)\mathbb{E}_{\{o,a\}\sim D_{\mathrm{online}},s\sim q_{\phi}}[\text{ELBO}].$ (4)
The amount of data we sample from the pretraining dataset is controlled by a hyper-parameter $\alpha$, which represents how much regularisation we put upon the model. In this paper, we mainly consider setting $\alpha=0.5$, so that we sample the data evenly from both datasets. We justify our choice in the ablation section with Figure 4(b).
This finding contradicts Rajeswar et al. [48] and Hansen et al. [23], where the pretrained world models do not need such a data-driven regulariser. We conjecture that unlike AIME they mainly use their world models purely as generative models to predict states and rewards given action sequences, which is only indirectly influenced by overfitting the ELBO.
3.2 Resolving the DKB
Based on the discussion from the previous sections, the straightforward way of solving the DKB is also to increase the number of demonstrations available to the agent. However, expert demonstrations are generally considered to be difficult and expensive to collect. Thus, increasing the size of the demonstration dataset is not always feasible in real-world applications.
In order to propose a more practical solution, we need to look deeper into what is the real cause of the DKB. The policy learning part of the AIME algorithm is essentially behaviour cloning, and it is only conducted on the demonstration dataset. So for the states covered in the demonstration dataset, the policy is given clear guidance about what to do, while for other states, the behaviour is undefined. AIME is solely relying on the generalisation abilities of the learned latent state and the trained policy network to extrapolate the correct behaviour. In particular for small demonstration datasets, this can be unreliable or even impossible. Therefore, if we can enlarge the space of the covered states, we should reduce the DKB [51].
Generative Adversarial Imitation Learning (GAIL) style algorithms [27, 45, 58, 37] are perfect examples of this solution. Instead of directly learning from the demonstration, they adversarially train a discriminator to assess how how closely each state resembles those in the demonstration dataset. When learning the policy, they treat the discriminator’s output as a reward and encourage visiting states that are more likely belonging to the demonstrations. This modification provides guidance for newly visited states in the replay buffer, allowing the space of supported states to grow over the course of training.
However, this benefit is not clearly separable in GAIL since it is always entangled with the adversarial training of the discriminator. A recent work MAHALO [35] shows further evidence of the importance of the size of the covered space. The authors studied a similar ILfO setup with embodiment and demonstration datasets. They compared two variants: for one they train an inverse dynamics model (IDM) from the embodiment dataset and use it to label the demonstration dataset, while for another they train a reward model from the demonstration dataset by labelling all time steps with a reward of $1$, and then use it to label the embodiment dataset. Finally, they run the same offline RL algorithm on both labelled datasets. The results show the second variant attains a much better performance even though the labelling from the reward model is not as meaningful as the actions from the IDM.
Based on these insights, we propose to introduce a surrogate reward providing guidance signal for the agent on the replay buffer dataset. Due to the instability of adversarial training [17, 2, 27] and our focus on the pretraining paradigm, we opt to adopt the VIPER algorithm [14]. Instead of training a discriminator, VIPER trains a video prediction model on the demonstration datasets and treats the likelihood of the video prediction model as the reward for policy learning, i.e.
$r^{\mathrm{VIPER}}_{t}=\log p_{\nu}(o_{t}|o_{<t}).$ (5)
Using this reward, we train the policy with a dreamer-style actor-critic algorithm [19] based on imagination in the latent space of the world model [19]. In order to do this, we first need to modify the reconstruction term in Equation 1 by adding an extra term for decoding the VIPER reward, i.e. $p_{\theta}(r^{\mathrm{VIPER}}_{t}|s_{t})$. Then, we further train a value estimator $V_{\xi}(s_{t})$ using TD($\lambda$)-return estimates, i.e.
$V^{\lambda}_{\xi}(s_{t})$ $=(1-\lambda)\sum_{n=1}^{\infty}{\lambda^{n-1}V^{(n)}_{\xi}(s_{t})}$ (6)
$\text{with}\;V^{(n)}_{\xi}(s_{t})$ $=\sum_{t^{\prime}=t+1}^{t+n}{\gamma^{t^{\prime}-t-1}r^{\mathrm{VIPER}}_{t^{\prime}}}+\gamma^{n}V_{\xi}(s_{t+n}).$
Using this estimate, we optimise our value function by minimising the MSE
$\xi^{*}=\mathop{\arg\!\min}_{\xi}(V_{\xi}(s_{t})-V^{\lambda}_{\xi^{\prime}}(s_{t}))^{2}.$ (7)
As is common practice, we use a target value network with parameters $\xi^{\prime}$ to stabilise training, whose parameters are updated using Polyak averaging with a learning rate $\tau$ in every iteration.
Using this value estimate, we extend the policy objective of Equation 3 to
$\psi^{*}=\mathop{\arg\!\max}_{\psi}\mathbb{E}_{o\sim D_{\mathrm{demo}},s\sim q_{\phi,\theta},a\sim\pi_{\psi}}[\text{ELBO}]+\beta\mathbb{E}_{\{o,a\}\sim D_{\mathrm{online}},s\sim q_{\phi},a^{\prime}\sim\pi_{\psi},s^{\prime}\sim p_{\theta}}[V_{\xi^{\prime}}^{\lambda}(s^{\prime})],$ (8)
where $\beta$ is a hyper-parameter for balancing the two terms. We set $\beta=0.1$ by default based on the difference of default learning rate in AIME and Dreamer.
4 Experiments
We aim to answer the following questions: a) How does the proposed AIME-v2 compare with state-of-the-art methods on common benchmarks? b) How well does the proposed modification resolve the EKB and the DKB? c) How do different choices of hyper-parameters influence the results? In order to answer these questions, we design our experiments on DMC and Meta-World benchmarks.
4.1 Datasets
For the DMC benchmark, we choose 9 tasks across 6 embodiments following Liu et al. [37] and use their published dataset as the demonstration datasets. Each dataset contains only 10 trajectories to reflect the scarcity of expert demonstrations. For the embodiment dataset, in order not to leak the task information from the pretraining phase, we follow Rajeswar et al. [48] and run a Plan2Explore [52] agent for each embodiment with 2M environments steps and use its replay buffer as the embodiment dataset. Different to them taking the model directly from the Plan2Explore agent as the pretrained model, we follow Zhang et al. [65] to retrain the model for $200\mathrm{k}$ gradient steps to get a better model.
For Meta-World benchmark, we use the data and model from Hansen et al. [23]. The embodiment dataset was created from the replay buffer datasets. The open-sourced replay buffer datasets contain $40\mathrm{k}$ trajectories for each of the $50$ tasks with only state information. In order to fit to our image observation setup, we render the images by resetting the environment to the initial state of each trajectory and then executing the action sequence. The details of resetting to the initial state can be found in Appendix D.
Following the idea of not leaking too much about the task information, inspired by the common practice in offline RL benchmarks [16], we use the first $200$ trajectories from each replay buffer and form a dataset with $10\mathrm{k}$ trajectories in total. We call this dataset MW-mt50. To further study the out-of-distribution transfer ability of the pretrained model, we follow the difficulty classification of the tasks from [53] and only use the $39$ easy and medium tasks to generate the datasets and the $11$ tasks hard and very hard tasks as hold-out tasks. We uniformly sample $250$ trajectories from te first $10\mathrm{k}$ trajectories from each of the $39$ tasks and form a dataset with $9750$ trajectories in total. We refer to this dataset as MW-mt39. Hence, MW-mt39 contains some expert behaviour solving the tasks, while MW-mt50 consists of mostly exploratory behaviour.
As the demonstration datasets, we use the single-task policies open-sourced by TD-MPC2 and collect $50$ trajectories for each tasks. We ensure that every trajectory in the demonstration dataset is successful. Since there are $500$ steps in a DMC trajectory and only $100$ steps in a Meta-World trajectory, the resulting datasets are roughly the same size.
Figure 2: Benchmark results and 9 DMC tasks. Return are calculated by running the policy 10 times with the environment and taking the average return. The results are averaged across 3 seeds with the shade region representing 95% CI.
4.2 Implementation
For the world model, we use the RSSM architecture [20] with the hyper-parameters in Hafner et al. [19] for DMC tasks. In addition, we use the KL Balancing trick from Hafner et al. [21] to make the training more stable. For Meta-World, since the visual scene is more complex, we use the M size model from Hafner et al. [22], but still with the continuous latent variable to be aligned with other models used in this paper. The policy network is implemented with a two-layer MLP, with $128$ neurons for each hidden layer. All the models are trained with Adam optimiser [31]. More details about the hyper-parameters can be found in Appendix C.
For the VIPER model, in the original paper, the authors first pretrain a VQ-GAN [15] and then train a GPT-style auto-regressive model in the quantised space for prediction. For simplicity of the implementation, in this paper, we consider training an unconditioned latent world model as in Seo et al. [54] to model the VIPER reward. We use the same RSSM architecture of the model learning for DMC, only removing the condition of the actions, and we train the VIPER model for each task separately. Especially during training, we find training such a powerful model from scratch on a small dataset can easily result in over-fitting. Thus, we empirically choose to train the model only for $500$ gradient steps for DMC models and $1000$ gradient steps for Meta-World models. We show evidence of overfitting in Appendix E. Due to the large scale of the ELBO, we also apply symlog [22] when computing the VIPER reward. Another difference with the original VIPER paper is that we do not use intrinsic motivation as the exploration bonus as the authors suggested, since the AIME loss for policy learning already provides task-related guidance for exploration. We only apply an entropy regulariser to the policy as is common practice. We further show the synergy between AIME and VIPER in Appendix F.
4.3 Benchmark Results
Figure 3: Benchmark results and $6$ Meta-World tasks. Trajectories are only counted as success when it success at the last time steps and the success rates are calculated with $10$ policy rollouts. The results are averaged across $3$ seeds with the shade region representing $95$% CI.
The benchmark results of DMC are shown in Figure 2. AIME-v2 significantly outperforms the PatchAIL baseline in $7$ out of $9$ tasks in terms of sample efficiency. Benefiting from the pretrained world model, AIME-v2 typically can reach expert performance within $200$k environment steps. Compared with BCO($\alpha$), updating the model is regularised and is benefitting more from the online interaction to resolve the EKB. Compared with the original offline AIME, AIME-v2 reliably improves performance, especially in hard tasks such as walker-run and hopper where offline AIME did not manage to make any progress.
However, there are still two tasks for which AIME-v2 does not show much progress, namely cartpole-swingup and quadruped-run. For cartpole-swingup, we observe that the policy learns to move the cart out of the scene so that the static image yields a high likelihood from the video prediction model. A similar phenomenon was also discussed in the original VIPER paper [14]. For quadruped-run, we conjecture that it is due to visual difficulties of a reconstruction-based model. When the quadruped is initialised on the ground, due to the symmetric structure of the robot, it is impossible to figure out which action corresponds to which leg, and it easily leads the action inference process to diverge. We additionally show AIME-v2 can work on these tasks with the help of the true reward in Appendix F.
The benchmark results of Meta-World are shown in Figure 3. We choose 4 tasks, namely disassemble, assembly, hand-insert and push from hard and very hard difficulties, and 2 tasks, namely sweep and hammer from the medium difficulty. While PatchAIL does not work on these tasks at all, AIME and AIME-v2 can make progress on them. AIME-v2 with either pretrained model outperforms AIME in all the tasks. On hard and very hard tasks, AIME with the mt50 model is better than AIME with the mt39 model. This is because they contain an unseen novel object by the mt39 model creating a large EKB. But in the online setting of AIME-v2, the two models are mostly on par. Moreover, using the mt50 models is better than using the mt39 models on average, which may imply covering diverse bahaviour is more valuable than knowing the expert directly.
4.4 Ablation Results
(a) Demonstration efficiency.
(b) Ablation of $\alpha$.
(c) Ablation of $\beta$.
Figure 4: (a) Performance of AIME-v2, MBBC, AIME with respect to different number of demonstrations. For AIME-v2, we do not show the result for more than $100$ demonstrations since it is already saturated to the expert. (b) Ablation for the choice of the regulariser ratio $\alpha$. The left figure shows the mean return over $10$ trajectories while the right figure shows the MSE between the inferred actions and the true actions. (c) Ablation for the choice of the weight of the value gradient loss $\beta$. All results are averaged across $3$ seeds with the shaded region representing a $95$% CI.
We conduct our ablation studies on walker-run task from DMC.
How well does the proposed methods resolve knowledge barriers? In order to show how well AIME-v2 resolves the two knowledge barriers, we conduct a similar experiment as Zhang et al. [65] by providing the agent with different numbers of demonstrations. The result is shown in Figure 4(a). As we discussed before, MBBC as an oracle method that circumvents the EKB is a strict upper bound for AIME. And AIME-v2 which addresses both the EKB and DKB achieves much better results and is an upper bound for MBBC. From AIME-v2 can achieve near-expert performance with as few as $5$ demonstrations for this challenging task.
Influence of the data regulariser ratio $\alpha$. We set the regulariser ratio $\alpha$ from $[0.0,0.25,0.5,0.75]$ and plot the results in Figure 4(b). As we can see from the result, as long as we enable the regulariser, i.e. set $\alpha>0$, we get reliable improvements over the course of training. But if we disable the regulariser by setting $\alpha=0$, the learning only starts once there is enough new data in the replay buffer. As we discussed in Section 3.1, without the regulariser, in the early stage of the training, the model can easily overfit to the replay buffer, and it explains the early flattening phase of the training. As the training progresses, more and more data is available from the replay buffer, and it can establish the regulariser on its own, which explains the dramatic growing phase of the curves.
We also plot the MSE between the inferred actions and the true actions during the training process. From that we can see that a higher regulariser ratio offers more stable inference of the actions in the early phase of training.
Influence of the value gradient loss weight $\beta$. We set the weight $\beta$ from $[0.0,0.01,0.05,0.1,0.5,1.0]$ and plot the results in Figure 4(c). As the result shows, having a small $\beta$ slows learning progress toward convergence. On the other hand, setting $\beta$ to a much larger value will improve the sample-efficiency without causing instability. For the sample efficiency, since we only have $10$ demonstrations, DKB dominates over EKB as shown in Figure 4(a). Thus, having a larger $\beta$ will make the learning much faster. In terms of the stability, as we discussed in 3.2, AIME loss and the value gradient loss operate on different regions of the environment states. This could make their influence on the policy independent of each other.
5 Related Work
Imitation Learning from Observation. ILfO [57, 58, 11, 35, 3, 65, 37] has become more popular in recent years due to their potential to utilise internet-scale videos for behaviour learning. Most of the previous works [57, 58, 35] study the problem only with the true state as observation. Recent works [11, 3, 65, 37] have started to shift toward image observations as a more general setting. Our work is a continuation of this journey.
Pretrained Models for Decision-Making. Inspired by the tremendous progress made in recent years in CV and NLP fields with the power of pretrained models, the decision-making community is also trying to follow the trend. Most recent works focus on the use of Large Language Model (LLM) for decision-making. A prompted model is used for producing trajectories and plans [10, 28, 1, 13], code [63, 36, 56, 9, 29] or for modifying the reward [39, 40]. There are also other people studying the benefit of pretrained visual models for visuomotor tasks [55, 41, 24, 44] while others try to train large policy networks directly with transformers [62] and huge datasets [8, 7, 49]. However, there is only little attention being put on pretrained world models [65, 48, 52], which are natively developed by the model-based decision-making community and perfectly fit into the pretraining and finetuning paradigm. Our work explores this overlooked domain and showcases its potential.
6 Discussion
In this paper, we identify two knowledge barriers, namely the EKB and the DKB, limiting the performance of state-of-the-art ILfO methods using pretrained models. We thoroughly analyse the underlying cause of each barrier and propose practical solutions. Specifically, we propose to use online interaction with a data-driven regulariser to solve the EKB and surrogate reward labelling to reduce the DKB. Combining these solutions, we propose AIME-v2, a method based on the recently proposed AIME and showcase its efficiency compared to state-of-the-art ILfO methods. We also conducted a series of ablation studies about how each knowledge barrier is addressed by the proposed solution and how their hyper-parameters influence the performance.
However, the proposed solutions still have drawbacks. First, the data-driven regulariser is not practical when the model is pretrained on huge datasets like the foundation models popular in the fields of CV and NLP. Reducing the amount of data needed for the regulariser could greatly improve the usability of the method. Second, although having pretrained models is beneficial, having too many pretrained components can be detrimental for model selection. Especially in AIME-v2, the world model and the VIPER model share a very similar interface. Designing a shared model that can serve both interfaces could ease the use of the method. Last but not least, due to the high demand of computing resources, we only study the pretrained world model on a very small scale. It will be an interesting direction to study these model at larger scales.
We hope our work can shed some light on the future development of ILfO method and bring more attention to the great potential of pretrained world models.
Appendix A Algortihm
Algorithm 1 AIME-v2
  Input: Embodiment dataset $D_{\mathrm{body}}$, Demonstration dataset $D_{\mathrm{demo}}$, Pretrained world model parameters $\phi,\theta$, Pretrained VIPER model $p_{\nu}$, regulariser ratio $\alpha$, value gradient weight $\beta$, batch size $\mathrm{B}$
  Initialise policy and critic parameters $\psi,\xi$ randomly.
  for $i=1$ to policy pretraining iterations do
     Draw a batch of demonstrations $o_{1:T}\sim D_{\mathrm{demo}}$
     Update policy parameters $\psi$ with Equation 3.
  end for
  Initialize $D_{\mathrm{online}}\rightarrow\emptyset$.
  for $i=1$ to Environment Interaction budget do
     Collect a new episode $\{o_{1:T},a_{1:T}\}$ with the current policy $\pi_{\psi}$
     Estimate reward using VIPER $r_{1:T}^{\mathrm{VIPER}}=p_{\nu}(o_{1:T})$
     Append $\{o_{1:T},a_{1:T},r_{1:T}^{\mathrm{VIPER}}\}$ to $D_{\mathrm{online}}$
     # Update world model
     Draw $\alpha\cdot\mathrm{B}$ samples $b_{\mathrm{body}}\sim D_{\mathrm{body}}$
     Draw $(1-\alpha)\cdot\mathrm{B}$ samples $b_{\mathrm{online}}\sim D_{\mathrm{online}}$
     Define combined batch $b=b_{\mathrm{body}}\cup b_{\mathrm{online}}$
     Finetune model with batch $b$ using Equation 4.
     # Update policy
     Sample a batch from $D_{\mathrm{demo}}$
     Update policy parameters $\psi$ with Equation 8.
     Update value function parameters $\xi$ with Equation 7.
  end for
Appendix B Compute Resources
All the experiments are run on a local cluster with a few A100 and RTX8000 instances. All the experiments are tuned to use less than 10GB of GPU memory so that they can run in A100 MIG. World models pretraining requires about 24 GPU hours, while VIPER models require negligible time for training. Each DMC experiment requires about 40 GPU hours while each Meta-World experiment requires about 24 GPU hours.
Appendix C Hyper-parameters
Here, we document the detailed hyper-parameters for all the trained models in Table 1.
Table 1: AIME-v2 hyper-parameters use for each benchmark.
DMC Meta-World
World Model
CNN structure Ha and Schmidhuber [18] Hafner et al. [22]
CNN width $32$ $48$
MLP hidden size $512$ $640$
MLP hidden layer $2$ $3$
MLP activations LayerNorm + Swish
Deterministic latent size $512$ $1024$
Stochastic latent size $30$
Free nats $1.0$
KL balancing $0.8$
Policy
Hidden size $128$
Hidden layer $2$
Activation ELU
Distribution Tanh-Gaussian
Value network
Hidden size $128$
Hidden layer $2$
Activation ELU
Target EMA decay $0.95$
Training
Batch Size $50$ $16$
Horizon $50$ $64$
Total Env steps $1\mathrm{M}$ $500\mathrm{k}$
Update ratio $0.1$
Gradient clip $100$
Policy entropy regulariser weight $1\mathrm{e}{-4}$
Model learning rate $3\mathrm{e}{-4}$
Policy learning rate $3\mathrm{e}{-4}$
Value network learning rate $8\mathrm{e}{-5}$
Discount factor $\gamma$ $0.99$
TD-Lambda parameter $\lambda$ $0.95$
Imagine horizon $15$
AIME-v2 specific
Policy pretraining iterations $2000$
Data-driven regulariser ratio $\alpha$ $0.5$
Value gradient loss weight $\beta$ $0.1$
Appendix D Details for Resetting Meta-World Tasks
To generate the image observation datasets from the TD-MPC2 replay buffer [23], we modify the Meta-World codebase to reset the environment to the initial state of the trajectory from the first observation. Luckily, the starting position of the robot arm is always the same for each task, so that we do not need to apply inverse kinematics to solve for the initial pose of the robot arm. For the object and the target position, for most of the tasks, the internal reset position can be computed by making a constant shift on the object position and the target position in the observations. There are, however, also a few edge cases which we handle differently.
In button-press-topdown and button-press-topdown-wall, the object’s true position only appears in the observation upon the second time step, presumably due to some simulator delay in the resetting process. So for these two tasks, the initial state is reset by the second observation.
For basketball and box-close, it seems like there is some internal collision detection that will alter the object and robot position after the task is reset, so computing the exact reset value from the observation is not possible. For these two tasks, we instead resort to a search-based method. To be specific, we use a gradient-free optimiser from [38] to search over the resetting space of the object and find the reset position that minimise the L2 distance with the true observation.
More details of the implementation can be found in the code.
Appendix E Overfitting of the VIPER Models
Figure 5: Correlation of the VIPER reward and the real reward with models trained with different number of gradient steps. Each point represents one trajectory. We can clearly see the model gradually overfitting and losing the correlation with the real reward when training for more than 1 00010001\,0001 000 gradient steps.
To better illustrate the overfitting problem for VIPER models and justify our choice of training fewer iterations, we train the VIPER models for a varying number of gradient steps and evaluate the correlations between the VIPER reward and the true reward on both the expert dataset from PatchAIL, where the VIPER model is trained on, and the replay buffer dataset from Zhang et al. [65].
Specifically, we train the same VIPER model with $\{100,500,1000,2000,5000,100000\}$ gradient steps and plot the result in Figure 5. As we can clearly see, when training with less than or equal to 1 00010001\,0001 000 gradient steps, VIPER reward has a very nice correlation with the true reward, with the middle-range performance even like a linear correlation. The best model could be selected from $500$ and $1000$ gradient steps. However, as we train the model for longer, the VIPER reward for the expert trajectories is boosted even higher, and as a side effect, it also relatively boosts up the VIPER reward for low-performance trajectories. This is because, when overfitting the expert trajectories, the model increases the marginal likelihood of all the observations in the expert trajectories to a higher value, which also includes a few frames of the robot lying on the ground at the very beginning of each trajectory after reset. For these low-performance trajectories, the robot remains mainly stuck around the initial position and struggles on the ground. This artifact of the overfitted VIPER reward creates a sharp local maximum in the low-performance region that the agent can hardly get away from.
Appendix F Additional Experiments
Figure 6: Additional ablation on DMC tasks by exploring the synergy between AIME and VIPER model. Return are calculated by running the policy 10 times with the environment and taking the average return. The results are averaged across 3 seeds with the shade region representing 95% CI.
Synergy between AIME and VIPER model. We also find there is a synergy between AIME and VIPER model. As we shown in Appendix E, one inherent problem of VIPER reward is that it not only incentivises the expert behaviour as the optimal, but also a stationary behaviour with very low reward as a local maxima. In order to work with the VIPER reward, the agent needs to have the ability to escape from the local maxima region. AIME offers the IL loss to imitate the expert demonstrations and can achieve decent performance even when pretrained offline, which helps to escape the local maxima. To better show the synergy, we provide additional ablation results with the VIPER reward in Figure 6. In the experiments, we include two other variants: the AIME-v2 w/o AIME is to remove the AIME IL loss from the online policy learning, so that the policy is pretrained by AIME loss but finetuned with only RL loss from the VIPER reward; while the VIPER is following the implementation in the original VIPER paper with RL loss on both the VIPER reward for the task and intrinsic reward for exploration. From the result, we can clearly see that without the help of AIME IL loss, VIPER reward cannot reliably motivate the agent to learn good behaviours. Even when in walker tasks, the w/o AIME variant can solve the tasks to certain extends, it depends strongly on the random seeds. In conclusion, the good performance of AIME-v2 cannot be achieved by either AIME IL loss or VIPER RL loss alone but a combination of both.
Figure 7: Additional benchmark results on 3 DMC tasks with an additional variant of AIME-v2 with the true reward. Return are calculated by running the policy 10 times with the environment and taking the average return. The results are averaged across 3 seeds with the shade region representing 95% CI.
Improving AIME-v2 with better rewards. We show additional results on the $3$ not so well-performing DMC tasks, namely cartpole-swingup, cheetah-run and quadruped-run, in 7. In the plot we add a new variant using the true reward from the environment to replace the VIPER reward. As the results show, if we had a better estimation of the surrogate reward, AIME-v2 could also achieve good performance on these tasks.